# Die Hard: The On-Off-Cycle of Galaxies on the Star Formation Main Sequence

### v1

@@ 6 Environmental Impact
It is well established that environment plays an important role for the star formation of satellites (Bahé et al. 2017; Lotz et al. 2019), but as we consider here only galaxies which remain as centrals until $z\approx 0$ the picture is less clear. In this case, Remus & Kimmig (2025) and Kimmig et al. (2025) do find indications that the environment plays a crucial role both for quenching as well as for rejuvenation, at least at high redshifts. This can be understood as star formation histories being dependent on both the availability of surrounding gas and the interaction with neighboring galaxies. A large number of lower-mass neighbors could fuel star-formation through minor mergers, while higher-mass neighbors could steal away ejected or loosely-bound gas and quench star formation in a similar fashion as for the Roche lobe of binary stars (e.g. Morris 1994).
Filaments have been proven as effective means to transport gas into the center while shielding against the quenching effect of a shock-heating hot halo (Dekel et al. 2009). While quenching can be induced by the intrinsic properties of a galaxy via hot halo and AGN, the frequent occurrence of rejuvenation processes suggests that the star formation history is at least in part determined by the environment.
In order to find systematic differences between the classes, we study two parameters through time:
  • $\bullet$
    The $d_{5}$ parameter quantifies the density of the environment, defined as the minimum radius at which five neighboring galaxies, each with a total mass greater than $10^{11}\ M_{\odot}$, are found (Remus & Kimmig 2025).
  • $\bullet$
@@
Figure 12: Violin distributions of the evolution of asphericities for each class at each snapshot, with main sequence rejuvenators and quenched rejuvenators combined as “Rejuvenators”. The asphericity of gas infall is evaluated on the spherical surface of each virial radius. The solid lines follow the median of each snapshot. The dotted lines are the median distributions of the other classes for a direct comparison. The main-sequence types maintain lower values and therefore a predominantly spherical gas infall. The other classes display peaking distributions between redshifts $0.52>z>$ 0.5.
@@
We find that galaxies which stay on the main sequence (cyan) also experience consistently isotropic infall throughout cosmic time, with only slightly higher asphericity between $0.52>z>$ 0.5. The other classes undergo a pronounced peak of aspherical infall, each at a different redshift. Caught-type galaxies are the first class to peak in asphericity at $z\approx 2$. Next are early-quenched galaxies which peak at $z\approx 1.2$. Last to peak in asphericity as a group are the late-quenched galaxies at $z\approx 0.7$. Rejuvenating galaxies peak at the same time as early-quenched galaxies, but with much lower total asphericity, indicating an infall history more similar to main-sequence-type galaxies.
This peaking distribution suggests a typical formation history of overall first isotropic infall, then a period of more filamentary accretion followed by a more isotropic flow that could reflect the buildup of halo fountains. Merger rates decline toward lower redshifts (O’Leary et al. 2021), such that increasing isotropy must be the result of smooth accretion. For the case of the main-sequence type galaxies, the rates are high enough (see Figure 13) that this gas is likely the result of increasing amounts of ejected gas inflowing back from all directions. We will study this process of fountain build-up in more detail in a future study.
@@ 7 Summary and Conclusions
We trace the evolution of $1282$ central galaxies in the highest-resolution box of the Magneticum Pathfinder cosmological hydrodynamical simulation suite through cosmic time, especially focusing on their gas inflow and star formation rates. By comparing their star formation rates with respect to expectations from star forming main-sequence observations by Pearson et al. (2018), we identify moments of sustained quenching and rejuvenating and classify the formation histories accordingly.
@@
  • $\bullet$
    Around $50\%$ (646) quench soon after the peak of star formation at $z\approx 2$, while $15\%$ (187) quench later between $01>z>$ 0.
  • $\bullet$
@@
This means that the dominant fraction of galaxies ($117$ of $159$) found on the star formation main sequence at $z\approx 0$ have undergone cycles of quenching and rejuvenation. These classes exist independent of which main sequence definition is used, for all main sequence definitions studied in this work (Speagle et al. 2014; Pearson et al. 2018; Leslie et al. 2020; Popesso et al. 2023), and only the relative frequencies per class change with definition.
These formation scenarios leave characteristic imprints on mean stellar ages and metallicities as well as the radial distribution of stellar ages, with rejuvenation generally taking place beyond the stellar half-mass radius. Galaxies that evolve along the main sequence consistently also have the most efficient star formation in their very centers, and the highest stellar metallicities.
@@
Acknowledgements
This work was performed at the LMU Munich with the support by the Excellence Cluster ORIGINS. The Magneticum simulations were performed at the Leibniz-Rechenzentrum with CPU time assigned to the Project pr83li. This work was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC-2094 - 390783311. We are especially grateful for the support by M. Petkova through the Computational Center for Particle and Astrophysics (C2PAP). LCK acknowledges support by the DFG project nr. 516355818. KD acknowledges support by the COMPLEX project from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program grant agreement ERC-2019-AdG 882679.
@@ Appendix A Varying the Smoothing Window
[…]
@@
Table 2: Numbers of galaxies within each formation class from top to bottom: main sequence, quenched, rejuvenating and caught with respect to expectations according to the star formation main sequence as found by Pearson et al. (2018). We compared the results using three smoothing windows of $0.5$, $1.0$ and $1.5$ gigayears each are compared in order to test the robustness of the classification.
@@
In order to check for robustness, we applied the algorithm using also a 0.5 and a 1.5 Gyr time window. The effect on sample sizes is shown in Table 2.
Larger smoothing windows disregard short-term fluctuations of star formation rates. As a consequence, more galaxies are considered main-sequence or quenched. Those would otherwise have been attributed to the group of rejuvenating galaxies, as short-term fluctuations would have been considered quenching and rejuvenating processes.
Larger smoothing windows leads to less registered rejuvenation cycles and therefore more galaxies falling into the main-sequence and quenched type galaxies; shorter windows have the opposite effect. The number of caught-type galaxies remains stable, as they have been classified manually. However, despite the varying class member numbers, each class roughly maintains its size, showing these characteristic evolution pathways emerge for a range of smoothing window size choices.
@@ Appendix B Gas Flows
We tested the link between gas flows and star formation by deriving the latter from flow rates and simplifying assumptions. The median net accretion rate evolution used for the derivation is shown in Figure 13. Specifically, we normalize each galaxy’s net accretion rate by respective expected accretion rate (Dekel et al. 2009) and then calculate class median values at each snapshot across cosmic time.
Figure 13: Median net gas accretion history evolution of each formation class across cosmic time. We show the net accretion rates at the virial radii as dashed lines and as solid lines accretion rates onto the center of the halos at ten percent virial radius. Prior to combining the data, we normalized each accretion rate value by halo mass and redshift according to expectations from dark-matter assembly (Dekel et al. 2009). The horizontal dashed black line at a value of one highlights this expectation by definition. The median of main-sequence galaxies are in cyan, quenched in dark red, rejuvenating galaxies in magenta and caught galaxies in blue colors. The vertical bars outline a representative 50th-percentile scatter. On average, accretion rates generally do not differ too much from the expectations. Quenching galaxies collectively assume negative values correlating with the time of quenching of their star formation. This correlation becomes evident from comparing the early- and late-quenching galaxy populations.
When comparing the median net accretion rates onto the galaxies at $0.1\ R_{\mathrm{vir}}$ (solid lines) for to the values at the virial radii (dashed lines) for each class, the figure shows that net gas accretion evolves similar at both radii. While main-sequence (cyan), caught (blue) and rejuvenator (magenta) type galaxies sustain positive net accretion rates roughly of the order of expectations from dark matter assembly (Dekel et al. 2009, dashed vertical line), quenching galaxies (dark red and dark blue) transition to even slightly negative values. Notably, the main-sequence type galaxies (cyan) have consistently higher values at the virial radii than at the galaxy borders.
@@
Appendix C Environment with the $d_{5}$ Parameter
@@
Figure 14: Evolution of the class median $d_{5}$ parameter values (Remus & Kimmig 2025) across cosmic time. Lower values indicate a denser environment of a galaxy. We chose comoving coordinates for a better comparison between across cosmic time. The 50th percentile scatter range shown in grey in the bottom left corner is similar for all classes and cosmic cosmic time. All classes show a very similar evolution with only faint differences with main-sequence types residing in denser environments consistently and caught-types found in less dense environments.

### v2

@@ 6 Environmental impact
It is well established that environment plays an important role for the star formation of satellites (bahe:2017; lotz:2019), but as we consider here only galaxies which remain as centrals until $z\approx 0$ the picture is less clear. In this case, remus:2023 and kimmig:2025a do find indications that the environment plays a crucial role both for quenching as well as for rejuvenation, at least at high redshifts. This can be understood as star formation histories being dependent on both the availability of surrounding gas and the interaction with neighboring galaxies. A large number of lower-mass neighbors could fuel star-formation through minor mergers, while higher-mass neighbors could steal away ejected or loosely-bound gas and quench star formation in a similar fashion as for the Roche lobe of binary stars (e.g. morris:1994).
Filaments have been proven as effective means to transport gas into the center while shielding against the quenching effect of a shock-heating hot halo (dekel:2009). While quenching can be induced by the intrinsic properties of a galaxy via hot halo and AGN, the frequent occurrence of rejuvenation processes suggests that the star formation history is at least in part determined by the environment.
seidel:2025 have shown that the Magneticum Pathfinder simulations are well-suited to study gas flows between halos and their environments, finding agreement in the flow-rates across both box resolution and size. While the gas mass resolution here is not high enough to trace the evolution of individual molecular clouds, we well resolve the more massive gas flows which are needed to rejuvenate galaxies in the mass ranges considered in this work. We therefore search for systematic differences between the environments of the classes with respect to the following two parameters:
    The $d_{5}$ parameter quantifies the density of the environment, defined as the minimum radius at which five neighboring galaxies, each with a total mass greater than $10^{11}\ M_{\odot}$, are found (remus:2023).
@@
Figure 12: Violin distributions of the evolution of asphericities for each class at each snapshot, with main sequence rejuvenators and quenched rejuvenators combined as “Rejuvenators”. The asphericity of gas infall is evaluated on the spherical surface of each virial radius. The solid lines follow the median of each snapshot. The dotted lines are the median distributions of the other classes for a direct comparison. The main-sequence types maintain lower values and therefore a predominantly spherical gas infall. The other classes display peaking distributions between redshifts $2>z>0.5$.
@@
We find that galaxies which stay on the main sequence (cyan) also experience consistently isotropic infall throughout cosmic time, with only slightly higher asphericity between $2>z>0.5$. The other classes undergo a pronounced peak of aspherical infall, each at a different redshift. Caught-type galaxies are the first class to peak in asphericity at $z\approx 2$. Next are early-quenched galaxies which peak at $z\approx 1.2$. Last to peak in asphericity as a group are the late-quenched galaxies at $z\approx 0.7$. Rejuvenating galaxies peak at the same time as early-quenched galaxies, but with much lower total asphericity, indicating an infall history more similar to main-sequence-type galaxies.
This peaking distribution suggests a typical formation history of overall first isotropic infall, then a period of more filamentary accretion followed by a more isotropic flow that could reflect the buildup of halo fountains. Merger rates decline toward lower redshifts (oleary:2021), such that increasing isotropy must be the result of smooth accretion. For the case of the main-sequence type galaxies, the rates are high enough (see Figure 13) that this gas is likely the result of increasing amounts of ejected gas inflowing back from all directions. We will study this process of fountain build-up in more detail in a future study.
@@ 7 Summary and conclusions
We trace the evolution of $1282$ central galaxies in the highest-resolution box of the Magneticum Pathfinder cosmological hydrodynamical simulation suite through cosmic time, especially focusing on their gas inflow and star formation rates. By comparing their star formation rates with respect to expectations from star forming main-sequence observations by pearson:2018, we identify moments of sustained quenching and rejuvenating and classify the formation histories accordingly.
@@
    Around $50\%$ (646) quench soon after the peak of star formation at $z\approx 2$, while $15\%$ (187) quench later between $1>z>0$.
@@
This means that the dominant fraction of galaxies ($117$ of $159$) found on the star formation main sequence at $z\approx 0$ have undergone cycles of quenching and rejuvenation. These classes exist independent of which main sequence definition is used, for all main sequence definitions studied in this work (speagle:2014; pearson:2018; leslie:2020; popesso:2023), and only the relative frequencies per class change with definition.
These formation scenarios leave characteristic imprints on mean stellar ages and metallicities as well as the radial distribution of stellar ages, with rejuvenation generally taking place beyond the stellar half-mass radius. Galaxies that evolve along the main sequence consistently also have the most efficient star formation in their very centers, and the highest stellar metallicities. The final black hole masses at $z\approx 0$ of main-sequence evolved galaxies and rejuvenating galaxies are distinctly lower than early-quenched galaxies at any given stellar mass, especially on the lower-mass end of the sample. Early formation conditions of efficient gas transport to the galaxies’ centers enhances not only star formation but also black hole growth, which in turn may impede rejuvenation or an evolution along the main sequence due to enhanced AGN feedback.
@@
Acknowledgements.
We thank the referee for their valuable feedback which helped with clarification and context of the results. This work was performed at the LMU Munich with the support by the Excellence Cluster ORIGINS. The Magneticum simulations were performed at the Leibniz-Rechenzentrum with CPU time assigned to the Project pr83li. This work was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC 2094 – 390783311. We are especially grateful for the support by M. Petkova through the Computational Center for Particle and Astrophysics (C2PAP). LCK acknowledges support by the DFG project nr. 516355818. KD acknowledges support by the COMPLEX project from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program grant agreement ERC-2019-AdG 882679.
@@ Appendix A Varying the smoothing window
Table 2: Numbers of galaxies within each formation class with respect to expectations according to the SFMS as found by pearson:2018.  [^3] From top to bottom: main sequence, quenched, rejuvenating and caught. We compared the results using three smoothing windows of $0.5$, $1.0$ and $1.5$ gigayears each are compared in order to test the robustness of the classification.
@@
In order to check for robustness, we applied the algorithm using also a 0.5 and a 1.5 Gyr time window. The effect on sample sizes is shown in footnote 3. Larger smoothing windows disregard short-term fluctuations of star formation rates. As a consequence, more galaxies are considered main-sequence or quenched. Those would otherwise have been attributed to the group of rejuvenating galaxies, as short-term fluctuations would have been considered quenching and rejuvenating processes. The number of caught-type galaxies remains stable, as they have been classified manually. However, despite the varying class member numbers, each class roughly maintains its size, showing these characteristic evolution pathways emerge for a range of smoothing window size choices.
@@ Appendix B Gas flows
We tested the link between gas flows and star formation by deriving the latter from flow rates and simplifying assumptions. The median net accretion rate evolution used for the derivation is shown in Figure 13. Specifically, we normalize each galaxy’s net accretion rate by respective expected accretion rate (dekel:2009) and then calculate class median values at each snapshot across cosmic time.
Figure 13: Median net gas accretion history evolution of each formation class across cosmic time. We show the net accretion rates at the virial radii as dashed lines and as solid lines accretion rates onto the center of the halos at ten percent virial radius. Prior to combining the data, we normalized each accretion rate value by halo mass and redshift according to expectations from dark-matter assembly (dekel:2009). The horizontal dashed black line at a value of one highlights this expectation by definition. The median of main-sequence galaxies are in cyan, quenched in dark red, rejuvenating galaxies in magenta and caught galaxies in blue colors. The vertical bars outline a representative 50th-percentile scatter. On average, accretion rates generally do not differ too much from the expectations. Quenching galaxies collectively assume negative values correlating with the time of quenching of their star formation. This correlation becomes evident from comparing the early- and late-quenching galaxy populations.
When comparing the median net accretion rates onto the galaxies at $0.1\ R_{\mathrm{vir}}$ (solid lines) for to the values at the virial radii (dashed lines) for each class, the figure shows that net gas accretion evolves similar at both radii. While main-sequence (cyan), caught (blue) and rejuvenator (magenta) type galaxies sustain positive net accretion rates roughly of the order of expectations from dark matter assembly (dekel:2009, dashed vertical line), quenching galaxies (dark red and dark blue) transition to even slightly negative values. Notably, the main-sequence type galaxies (cyan) have consistently higher values at the virial radii than at the galaxy borders.
@@
Appendix C Influence of AGN feedback
piotrowska:2022 have shown that AGN feedback quenching is strongly related to the mass of supermassive black holes in the centers of galaxies. We therefore show in Figure 14 the black hole masses and stellar masses of each galaxy at $z\approx 0$ colored according to its star formation evolution class to estimate the influence of AGN feedback on the evolution scenarios. At any given stellar mass, quenched galaxies (dark red and dark blue circles) occupy the highest black hole masses. This is likely due to how gas flows onto the central black hole, with high-z environments creating more isotropic flows (kimmig:2025a, Seidel et al. in prep), while galaxies which make a majority of their stars at later times or in on-off cycles do not as efficiently funnel the gas toward the center. This leads to lower black hole masses, with the most extreme case found for those evolving along the main sequence. We posit that galaxies which solely grew along the main sequence are the result of higher angular momentum gas that forms its stars in a disk without strongly feeding the central supermassive black hole.
Figure 14: $\mathrm{M}_{\mathrm{BH}}$-$\mathrm{M}_{*}$ relation of the galaxy sample at $z\approx 0$. Especially for the mass range of $10^{10}\mathrm{M}_{\odot}\leq\mathrm{M}_{\odot}\leq 10^{11}\mathrm{M}_{\odot}$, a clear transition order from quenched galaxies along the Magorrian relation to lower black hole masses for star forming galaxies. Specifically, early-quenched galaxies (dark red) have the highest values, followed by late-quenched (dark blue), rejuvenated quenched (lilac), MS-evolved (cyan) and finally rejuvenating MS (pink) galaxies as well as caught-type galaxies (blue) at the lower black hole mass end.
Appendix D Environment with the $d_{5}$ parameter
@@
Figure 15: Evolution of the class median $d_{5}$ parameter values (remus:2023) across cosmic time. Lower values indicate a denser environment of a galaxy. We chose comoving coordinates for a better comparison between across cosmic time. The 50th percentile scatter range shown in gray in the bottom left corner is similar for all classes and cosmic cosmic time. All classes show a very similar evolution with only faint differences with main-sequence types residing in denser environments consistently and caught-types found in less dense environments.